OpenStudy (anonymous):

help me with the 3rd one

OpenStudy (alexandervonhumboldt2):

yes

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

use ratio: 360/40=x/60

OpenStudy (alexandervonhumboldt2):

x=360/40*60=9*60=?????

OpenStudy (anonymous):

c

OpenStudy (alexandervonhumboldt2):

nope

OpenStudy (anonymous):

d

OpenStudy (alexandervonhumboldt2):

yes